# Controllable Learning in Information Retrieval: A Survey

### v1

@@ 3.1 Who Controls
“Who controls” refers to who proposes the task description $\bm{s}_{\mathrm{desc}}$, whether it is the user or the platform. When users have different needs, such as wanting to protect their privacy, or the platform wishes to increase result diversity to ensure different items get exposure, a corresponding task description arises. This section will introduce existing work based on who proposes the task description to control the IR systems.
@@ 3.1.1 User-Centric Control
Systems like UCRS [1], IFRQE [2], and LACE [3] empower users to actively shape their recommendation experiences. UCRS allows users to issue control commands to mitigate filter bubbles [1], IFRQE lets users decide which of their interactions should contribute to the training of the recommendation model [2], and LACE enables users to edit concept-based profiles for personalized text recommendations [3]. These systems exemplify user-centric control by providing users with the tools to adjust and refine the recommendation process according to their preferences and needs. Additionally, the user-controlled recommender system for connecting students with advisors further emphasizes user-centric control by leveraging user feedback and selections to guide the recommendation process, ensuring that the recommendations are aligned with the user’s academic or research goals [4].
@@ 3.1.2 Platform-Mediated Control
Platform-mediated control is characterized by algorithmic adjustments and policy-based constraints that a platform can impose on the recommendation process. For instance, recent work aims to consider a general platform-side framework to control the algorithms in recommender systems and to avoid the creation of isolated communities and the filter bubble effect [8]. ComiRec captures multiple interests from user behavior sequences and generates the top N items controllable by the platform in the aggregation module [5]. The CMR framework employs a hypernetwork to dynamically generate parameters for re-ranking models based on different preference weights, allowing the platform to adapt recommendations to various user groups or environmental changes without requiring user intervention [6]. CCDF platform controls the diversity of categories in recommendations by selecting top categories based on user-category matching and constrained item matching, aiming to alleviate echo chamber effects and broaden user interests [7]. Huan et al. [57] introduces a groundbreaking approach to addressing the complexities of heterogeneity in MSR. By deploying the Scenario Aware Model-Agnostic Meta Distillation (SAMD) framework, the research offers a method for scenario-aware and model-agnostic knowledge dissemination across diverse contexts, enhancing the system’s ability to tailor recommendations precisely. Some studies introduce specialized networks to meet the controllable demands. DTRN [58] obtains task-specific bottom representation explicitly by making each task have its own representation learning network in the bottom representation modeling stage. HyperBandit [9] maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. These systems demonstrate how platform-mediated control can optimize recommendation strategies to meet specific objectives, such as increasing diversity or improving conversion rates, while maintaining the quality of recommendations.
3.2 What is Controllable
This section aims to clarify “what is controllable” for a controllable learner in information retrieval applications. Specifically, we explore the question: “What is task target $\bm{s}_{\mathrm{tgt}}$ in Definition 1 of controllable learning?” A task target emerges when preferences shift across multiple objectives, such as accuracy and diversity. Another task target involves ensuring that learners can adapt to their environment, given descriptions of various environmental elements. Additionally, the context $\bm{s}_{\mathrm{ctx}}$ also needs to be controllable. For example, certain sensitive information in $\bm{s}_{\mathrm{ctx}}$ needs editing, while the target requires ensuring that the controllable learner’s outputs remain consistent with a model trained based on the edited contexts. Following this classification, this section reviews existing research to identify which target can be effectively achieved.
3.2.1 Multi-Objective Control
Objectives like precision, recall, and diversity, are at the core of recommender system performance metrics. ComiRec [5] strikes a balance between accuracy and diversity through hyperparameters in the aggregation module by capturing multiple user interests and integrating them into the recommendation process. UCRS [1] introduces a system where users can proactively reduce filter bubbles, controlling the recommendations they receive via hyperparameters that affect accuracy and coverage. The CMR framework [6] uses a hypernetwork to dynamically adjust preference weights, aiming to balance accuracy, diversity, etc. through platform-defined preference weights. These systems collectively focus on achieving a harmonious balance among the various objectives of recommendation quality, ensuring that users receive accurate, diverse, and relevant recommendations.
3.2.2 Historical Behavior Control
Some work in recommender systems involves allowing users to manage the context including their past interactions and profile information. LACE [3] allows users to edit their profiles by selecting or deselecting concepts, directly influencing the recommendations. LFRQE [2] lets users delete or modify past interactions to refine their recommendation experience. Grapevine [4] makes personalization more transparent and user-controllable using “open user models”, making normally hidden user models visible to end users and allowing them to change their values, thereby affecting the results of personalization. Together, these methods empower users to shape their digital footprint and ensure that recommendations better align with their evolving interests and preferences.
3.2.3 Controllable Environmental Adaptation
Compared to dynamic online learning [59, 60, 61], Controllable environmental adaptation is an explicit method that involves describing environmental elements and inputting them into the controllable learner, rather than solely relying on environmental feedback to update the model. HyperBandit [9] utilizes a hypernetwork to explicitly model the binary relationship between periodic external environments and user interests. In the testing stage, HyperBandit takes time features as input and dynamically adjusts the recommendation model parameters to account for time-varying user preferences.
3.2.4 Other Targets
In addition to the aforementioned control targets, there are many other forms of task targets, such as Multi-Scenario Control[58] and Multi-Domain Control[24], depending on the application context of the CL algorithms.
@@ 4 How Control is Implemented
In Section 3, the existing works are categorized based on controllers and controllable objectives. In this chapter, we summarize the common control methods and their application positions), corresponding to the CL techniques and application position (i.e., where to control). Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “CL tech.” and “where” categories (see  Table I).
@@ 4.1 Controllable Learning Techniques
We analyze and summarize the CL techniques for implementation of the control function $h$, where $h$ maps the learner $f$ to a new learner $f_{\mathcal{T}}$ based on the triplet $\mathcal{T}$.
@@ 4.1.1 Rule-Based Techniques
Rule-based controllable techniques have proven to be indispensable in many scenarios such as recommender systems, information retrieval, and text generation due to its direct series mode after the upstream models. These methods, acting as a patchwork solution, enhance system performance such as security, fairness, and interpretability from a product perspective. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents the system performance we expect to achieve, such as security, fairness etc. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ is always implicit, embedded in the design rationale of pre-processing methods like user profiles modification and post-processing methods, such as the removal of offensive language in text generation and enhancing the diversity in information retrieval. Conversely, $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ always represents the output from upstream models, including recommendation lists, retrieved documents, and raw generated texts.
@@
Rule-based techniques in information retrieval systems involves applying predefined rules to the outputs of AI models to refine and enhance the results. Parikh [10] discusses the use of augmented AI and computer vision to extract information from unstructured data. By employing rule-based methods, the system can ensure that the extracted information is accurate and relevant, improving the overall quality and reliability of the information retrieval process. Khosrobeigi et al. [11] present a rule-based approach to improve the performance of Persian Optical Character Recognition (OCR). By incorporating rule-based corrections, the system can significantly reduce errors and enhance the readability and correctness of the output text. Kang et al. [12] demonstrate the application of rule-based natural language processing (NLP) techniques to improve disease normalization in biomedical texts. This approach ensures that the generated text not only meets the syntactic requirements but also maintains semantic consistency and domain accuracy. Recommender systems also benefit greatly from rule-based post-processing to ensure fairness, diversity, and explainability of recommendations. Nandy et al. [13] explore how fairness can be achieved through post-processing in web-scale recommender systems. By applying rule-based adjustments, the system can mitigate biases and ensure that recommendations are fair and equitable across different user groups. Le et al. [14] combine embedding-based and semantic-based models for post-hoc explanations in recommender systems. Rule-based post-processing plays a crucial role in providing clear and understandable explanations for the recommendations, enhancing user trust and system transparency. Antikacioglu and Ravi  [15] focus on improving diversity in recommender systems through post-processing techniques. By incorporating rules that promote diversity, the system can offer a broader range of recommendations, preventing the issue of over-specialization and improving user satisfaction.
@@ 4.1.2 Pareto Optimization
In the domain of machine learning, the concept of controllability has garnered significant attention due to the necessity to balance multiple, often conflicting objectives inherent in lots kinds of tasks. The integration of Pareto optimality principles from multi-objective optimization (MOO) has proven particularly instrumental in crafting algorithms that can navigate these trade-offs effectively. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents a single Pareto optimal solution that simultaneously satisfies multiple target objectives. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ generally consists of a set of multi-objective weights or constraints that enable the system to achieve the target. The $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context.
@@
Lin et al. [62] pioneers this field by proposing an algorithm that decomposes the MTL problem into a series of subproblems, each representing a unique trade-off preference. This approach yields a set of Pareto optimal solutions, allowing for flexibility in selecting the most appropriate solution based on specific task requirements. Furthermore, Lin et al. [19] advances this concept by introducing a framework that enables real-time adjustments to these trade-offs, thereby providing a more dynamic and adaptable system capable of responding to changing user preferences and task demands. In parallel, the quest for scalable solutions has led to the exploration of methods that approximate the Pareto front, a representation of the optimal trade-offs among objectives. Ruchte et al. [18] propose a technique that conditions deep neural networks directly on preference vectors, thereby generating a well-distributed set of Pareto solutions. This work addresses the computational challenges associated with deep learning and sets the stage for more efficient and effective exploration of the Pareto front in complex MTL scenarios. A novel framework PHN-HVI [16] is proposed, which uses a hypernetwork to generate diverse solutions and improve the quality of the Pareto front in multi-objective optimization problems by optimizing the Hypervolume indicator, significantly outperforming existing methods. Pareto HyperNetworks (PHNs) [17] is introduced by Navon et al., where a single hypernetwork approach is applied to efficiently learn the entire Pareto front for multi-objective optimization problems, generalizes to unseen operating points, and enables post-training selection of optimal models based on runtime preferences. Complementing these efforts, research has also focused on the continuous exploration of Pareto sets, which provides a more nuanced understanding of the trade-offs in MTL. Ma et al. [20] introduce an innovative method that constructs continuous approximations of local Pareto sets, facilitating the applicability of MOO principles to the realm of deep MTL. Some researches have encompassed multi-objective optimization and Pareto efficiency in recommendation algorithms [21][22] [23]. Lin et al. [21] introduces a general framework for reconciling objectives such as Gross Merchandise Volume (GMV) and Click-Through Rate (CTR) in E-Commerce Recommendation. This work is complemented by Ribeiro et al. [22], which proposes novel methods for aggregating and hybridizing existing recommendation algorithms to achieve a Pareto-efficient ranking. Similarly, a hybridization method for multi-objective in recommender systems is proposed by Ribeir et al. [23], which contributes to this discourse by employing an evolutionary search to combine algorithms that excel in different objectives, leading to a Pareto frontier that can be adjusted according to user needs.
@@ 4.1.3 Hypernetwork
In addition to implicit control methods such as retraining and fine-tuning used to regulate models, hypernetworks have gradually emerged as a key technique for enhancing model controllability due to their explicit control capabilities. The advent of hypernetworks has revolutionized the landscape of artificial intelligence, particularly in the realm of parameter control for various tasks such as information retrieval and recommender systems. Hypernetworks, which are neural networks designed to generate the parameters for another network, offer a flexible and efficient way to manage and adapt model parameters dynamically. Following Definition 1, the $s_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents an expected result output by the controllable learner in the specific domain or task. The $s_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ typically represents the description of the task or domain, usually serving as the input for the hypernetwork. The hypernetwork outputs the parameters for the controllable learner, enabling it to achieve the corresponding $s_{\mathrm{tgt}}$. The $s_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context.
@@
Galanti and Wolf [27] discuss the modularity of hypernetworks, emphasizing their ability to decompose a complex task into smaller, manageable sub-tasks. This modular approach allows for the generation of highly specialized parameters tailored to specific retrieval tasks, thereby improving accuracy and efficiency. He et al. [25] introduce Hyperprompt, a prompt-based task-conditioning method for transformers that leverages hypernetworks. This approach dynamically generates prompts that condition the transformer model to adapt to different retrieval tasks, enabling more precise and context-aware information retrieval. Li et al. [24] propose Hamur, a hyper adapter designed for multi-domain recommendation. Hamur utilizes hypernetworks to adaptively generate domain-specific parameters, thereby enhancing the model’s ability to recommend items across different domains with high relevance and accuracy. Furthermore, Chen et al. [6] explore the use of policy hypernetworks for controllable multi-objective re-ranking in recommender systems. By leveraging hypernetworks to generate parameters that balance multiple objectives, such as relevance and diversity, the system can dynamically adjust its recommendations based on user preferences and system goals. Similarly, Shen et al. [9] propose HyperBandit, whcih maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Mahabadi et al. [26] discuss parameter-efficient multi-task fine-tuning for transformers via shared hypernetworks. This method allows for the efficient sharing of parameters across multiple tasks, reducing the overall computational load and improving the scalability of transformer models. The shared hypernetwork generates task-specific parameters that fine-tune the model for each particular task, enhancing its adaptability and performance. In conclusion, hypernetworks represent a powerful tool for parameter control in information retrieval and recommender systems. By enabling dynamic and task-specific parameter generation, hypernetworks can significantly improve the performance and adaptability of AI models, paving the way for more sophisticated and responsive systems.

### v2

@@ 3.2 Who Controls
This section aims to clarify “who controls”, which refers to who proposes the task description $\bm{s}_{\mathrm{desc}}$ in Definition 2.1. According to our research, controllability in recommendation systems can be categorized from the perspective of “who controls” into two aspects: user-centric and platform-centric. Users may have various needs, such as protecting their privacy, removing accidental interactions from their history (i.e., data filtering), indirectly expressing preferences [75], and explore the novel items. The platform may aim to increase the diversity of the learner’s output to promote the exposure of less popular items, adjust the trade-off between multi objectives (e.g., accuracy and diversity), utilize the adaption between multi scenarios to promote the performance, and unify controllable model for all to cut the cost (i.e., efficiency). The objectives of each side are shown in Figure 6.
Figure 6: The Objects of User-Centric Control and Platform-Centric Control.
@@ 3.2.1 User-Centric Control
User-Centric Control focuses on explicitly defining users’ preferences for specific content. This means that users have a clear awareness of their own interests (i.e., task target) and provide this information to the recommendation system in a specified format (i.e., task description), such as through questionnaires, weighting buttons, or natural language. Additionally, users may not explicitly express their preferences. Instead, they may convey them through behaviors, such as interactions with recommendation results.
Explicit Control. In some applications, users explicitly defining their preferences, which means that users have a clear awareness of their own interests (i.e., task target). Then the users clearly provide this information to the recommendation system in a specified format (i.e., task description), such as through questionnaires, weighting buttons, or natural language.
Supervised $\beta$-VAE [15] proposes a supervised disentangled $\beta$-VAE-based recommendation model, which explicitly assigns latent dimensions to user preference attributes and aligns them with labeled features to achieve disentanglement. By incorporating classification loss for effective supervision on partially labeled data, the approach generates interpretable and controllable recommendations while maintaining high recommendation performance and enabling fine-grained user control over the output. LP [16] presents a method allowing users to explicitly control recommendation results through iterative critiques. It extends a linear embedding-based recommendation approach by integrating subjective keyphrase feedback and modulating critique strength in a multistep process. This method reduces interactions required for satisfactory recommendations and improves success rates, showcasing enhanced user-centric recommendation adaptability. TEARS [9] captures user preferences in a text-based format and leverages optimal transport to align embeddings from summaries with black-box embeddings based on user history during training. This setup allows users to edit their summaries directly, giving them control over the recommendations they receive during testing. LangPTune [17], a framework that enables explicit user control over recommendation systems through natural language profiles. By integrating end-to-end training with reinforcement learning and contrastive learning, it optimizes both recommendation accuracy and the interpretability of user preferences. This approach ensures users can directly modify their profiles to steer recommendations, enhancing transparency and personalization.
Implicit Control. Users may not explicitly express their preferences. Instead, they may convey them through behaviors, such as interactions with recommendation results. Under this setting, the task description may not exist. The user’s preference can be reflected through modifications to the context, which encompasses interaction history, general user profile and item description.
UCRS [2] proposes four control commands to reduce filter bubbles. Fine-grained user control allows users to increase items associated with their specific characteristics. Coarse-grained user control helps users reduce items tied to their own profile traits. Fine-grained item control enables adjustments like increasing items from specific categories. Coarse-grained item control allows reducing items from the most frequent category in the user’s interaction history. In practical scenarios, users may not want all their behaviors to be included in the model training process. IFRQE [7] lets users decide which of their interactions should contribute to the training of the recommendation. The models are then optimized to maximize utility, which considers the trade-off between recommendation performance and respecting user preferences. LACE [6] enables users to edit concept-based profiles for personalized text recommendations. LACE represents each user as a set of human-readable concepts by retrieving items the user has interacted with. These concept-based and editable user profiles are then used to make recommendations. The model’s design enables intuitive interaction with transparent user profiles to control recommendations.
However, whether explicit or implicit, these distinctions merely represent a form of user-controllable implementation. Their underlying purpose remains consistent: to foster interest, protect privacy, eliminate noise, and explore the unknown.
@@ 3.2.2 Platform-Mediated Control
Platform-mediated control is characterized by algorithmic adjustments and policy-based constraints that a platform can impose on the recommendation process. In terms of form, control requirements are still expressed as $\bm{s}_{\mathrm{desc}}$. However, unlike user-centric control, platform-mediated control represented by $\bm{s}_{\mathrm{tgt}}$ focuses more on the adjustment to various and dynamic tasks during the test time phase.
@@
For instance, recent work aims to consider a general platform-side framework to control the algorithms in recommender systems and to avoid the creation of isolated communities and the filter bubble effect [19]. Coincidentally, CCDF [18] controls the diversity of categories in recommendations by selecting top categories based on user-category matching and constrained item matching, aiming to alleviate echo chamber effects and broaden user interests. Introducing diversity alone may lead to a decline in recommendation performance, prompting the platform to focus on controlling multiple objectives. The CMR [3], Padirec [4] both utilize parameter generation for downstream models based on different objective weights, allowing the platform to adapt recommendations to various user groups or environmental changes at test time. In addition to diversity, user interests are also multifaceted. ComiRec [1] captures multiple interests from user behavior sequences and generates the top N items controllable by the platform in the aggregation module. Additionally, scenario signals and task signals are also utilized as control conditions to enhance the model’s adaptability. SAMD [20] offers a method for scenario-aware and model-agnostic knowledge dissemination across diverse contexts, enhancing the system’s ability to tailor recommendations precisely. HyperBandit [12] takes into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Additionally, DTRN [21] obtains task-specific bottom representation explicitly by making each task have its own representation learning network in the bottom representation modeling stage.
In summary, the platform-mediated control primarily aims to optimize recommendation strategies to meet specific objectives, such as increasing diversity or improving recommendation accuracy. Beyond these goals, there are more complex control tasks, such as control among multiple objectives and the transitions between scenarios and tasks. Overall, platform-mediated control focuses on equipping the model with the ability to adapt to specific task directives during the test time phase.
@@ 4 How Control is Implemented
In Section 3, the existing works are categorized based on controllable objectives (i.e.,what) and controllers (i.e., who). In this section, we summarize the common control methods and their application positions within the recommendation system workflow, corresponding to Section 4.1 and Section 4.2.
@@ 4.1 Controllable Learning Techniques
We analyze and summarize the controllble learning (CL) techniques for implementation of the control function $h(\cdot)$, where $h(\cdot)$ maps the learner $f$ to a new learner $f_{\mathcal{T}}$ based on the triplet $\mathcal{T}$ following the definition in Definition 2.1. The control function is regarded as a key mechanism for exerting control over the learner. However, to the best of our knowledge, there are several variations of this framework, which we will introduce in the following sections.
@@ 4.1.1 Rule-Based Techniques
Rule-based controllable techniques have proven to be indispensable in many scenarios such as recommender systems, information retrieval, and text generation due to its direct mode. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ of the triplet typically represents the system performance we expect to achieve, such as security, fairness etc and the $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ is always implicit. Rule-based methods acting as a patchwork solution, enhance system performance from a product perspective. Specifically, rule-based methods can be categorized into two types: pre-processing and post-processing. Formally, the control function $h$ is defined as a mapping: $h:f\rightarrow f\circ g_{\text{rule}}$ in the context of pre-processing, whereas it is defined as: $h:f\rightarrow g_{\text{rule}}\circ f$ in the context of post-processing, where $f$ represents the base function and $g_{\text{rule}}$ denotes the rule-based control mechanism. For pre-processing algorithms, the strategy involves handling the context information including user profiles, interaction history, and others, denoted as $\bm{s}_{\mathrm{ctx}}$ in Definition 2.1, with the aim of altering the representations of users and items to meet specific requirements, such as privacy preservation or fairness. In contrast, post-processing algorithms, such as removing outdated items or promoting less popular items, can directly enhance diversity.
@@
Rule-based techniques involves applying predefined rules to the inputs or outputs of recommender systems to refine and enhance the data or results. In terms of privacy protection, the platform can achieve security and control through modifications and synthesis of input data [76, 77, 78]. In the aspect of diversity, rule-based heuristic algorithms for recommended items based on user needs can also prove effective [39, 79]. Specifically, MMR [39] iteratively selects items that maximize a combined criterion of relevance to the user’s query and minimal redundancy with already selected items, which ensures that the final recommendation list is both pertinent and varied. Additionally, rule-based methods are also highly effective in personalization and fairness. Nandy et al. [25] explore how fairness can be achieved through post-processing in web-scale recommender systems. By applying rule-based adjustments, the system can mitigate biases and ensure that recommendations are fair and equitable across different user groups. Similarly, Antikacioglu and Ravi  [27] focus on improving diversity in recommender systems through post-processing techniques. By incorporating rules that promote diversity, the system can offer a broader range of recommendations, preventing the issue of over-specialization and improving user satisfaction. In addition, in the explainability of recommendations, rule-based methods have demonstrated a simple yet effective performance. Le et al. [26] combine embedding-based and semantic-based models for post-hoc explanations in recommender systems. Rule-based post-processing plays a crucial role in providing clear and understandable explanations for the recommendations, enhancing user trust and system transparency.
@@ 4.1.2 Pareto Optimization
In the domain of machine learning, the concept of controllability has garnered significant attention due to the necessity to balance multiple, often conflicting objectives inherent in lots kinds of tasks. The integration of Pareto optimality principles from multi-objective optimization (MOO) has proven particularly instrumental in crafting algorithms that can navigate these trade-offs effectively. Following Definition 2.1, the $\bm{s}_{\mathrm{tgt}}$ of the triplet typically represents a single Pareto optimal solution that simultaneously satisfies multiple target objectives. The $\bm{s}_{\mathrm{desc}}$ generally consists of a set of multi-objective weights or constraints that enable the system to achieve the target. The $\bm{s}_{\mathrm{ctx}}$ usually depends on the scenario. In the field of information retrieval, it often includes candidate documents, user profiles, and other relevant context. As illustrated in Figure 7, the key challenge of Pareto optimization in the domain of controllable learning is determining how to guide the learner $f$ to achieve Pareto optimality under given constraints.
@@
Pareto MTL [80] pioneers this field by proposing an algorithm that decomposes the MTL problem into a series of subproblems, each representing a unique trade-off preference. This approach yields a set of Pareto optimal solutions, allowing for flexibility in selecting the most appropriate solution based on specific task requirements. Furthermore, Controllable Pareto MTL [31] advances this concept by introducing a framework that enables real-time adjustments to these trade-offs, providing a more dynamic and adaptable system capable of responding to changing user preferences and task demands. In parallel, research on user-side personalized preferences has also emerged progressively. Ruchte et al. [30] propose a technique that conditions deep neural networks directly on preference vectors, thereby generating a well-distributed set of Pareto solutions. This work addresses the computational challenges associated with deep learning and sets the stage for more efficient and effective exploration of the Pareto front in complex MTL scenarios. A novel framework PHN-HVI [28] is proposed, which uses a hypernetwork to generate diverse solutions and improve the quality of the Pareto front in multi-objective optimization problems by optimizing the Hypervolume indicator, significantly outperforming existing methods. Similarly, PHNs [29] utilizes a single hypernetwork approach to efficiently learn the entire Pareto front for multi-objective optimization problems, generalizes to unseen operating points, and enables post-training selection of optimal models based on runtime preferences. Complementing these efforts, research has also focused on the continuous exploration of Pareto sets, which provides a more nuanced understanding of the trade-offs in MTL. Ma et al. [32] introduce an innovative method that constructs continuous approximations of local Pareto sets, facilitating the applicability of MOO principles to the realm of deep MTL. Some researches have encompassed multi-objective optimization and Pareto efficiency in recommendation algorithms [81, 33]. PAPERec [81] introduces personalized Pareto optimality in multi-objective recommendation achieving adaptive control of personalized objective weights. This model allows for dynamic balancing of conflicting objectives, such as accuracy and diversity, tailored to individual user preferences. Lin et al. [33] introduces a general framework for reconciling objectives such as Gross Merchandise Volume (GMV) and Click-Through Rate (CTR) in E-Commerce Recommendation. MoFIR [82] leverages multi-objective reinforcement learning to provide controllability over the trade-off between utility and fairness in recommendation systems. By introducing a conditional network and a Pareto-efficient optimization approach, the framework allows decision-makers to dynamically adjust recommendation outcomes based on predefined utility-fairness preferences.
Most of the aforementioned studies aim to achieve customized Pareto optimality. However, they do not focus on test-time control, which means the learner $f$ still requires retraining to accommodate new user requirements and achieve Pareto optimality. We believe that test-time control to achieve Pareto optimality without retraining is a promising research direction and anticipate increased attention to this area in the future.
@@ 4.1.3 Hypernetwork
Hypernetworks [83], which are neural networks designed to generate the parameters for another network, offer a flexible and efficient way to manage and adapt model parameters dynamically. Therefore, they have gradually emerged as a key technique for enhancing model controllability due to their explicit control capabilities. Following Definition 2.1, the $\bm{s}_{\mathrm{desc}}$ typically represents the description of the task or domain, usually serving as the input for the hypernetwork. The hypernetwork act as the control function $h$ taking $\bm{s}_{\mathrm{desc}}$ as input, mapping the learner through parameter generation. Thus the mapped learner could achieve the corresponding $\bm{s}_{\mathrm{tgt}}$.
@@
Galanti and Wolf [37] discuss the modularity of hypernetworks, emphasizing their ability to decompose a complex task into smaller, manageable sub-tasks. This modular approach allows for the generation of highly specialized parameters tailored to specific retrieval tasks, thereby improving accuracy and efficiency. Hyperprompt [36] introduces a prompt-based task-conditioning method for transformers that leverages hypernetworks. This approach dynamically generates prompts that condition the transformer model to adapt to different retrieval tasks, enabling more precise and context-aware information retrieval. Hamur [13] propose a hyper adapter designed for multi-domain recommendation. Hamur utilizes hypernetworks to adaptively generate domain-specific parameters, thereby enhancing the model’s ability to recommend items across different domains with high relevance and accuracy. Furthermore, CMR [3] explore the use of policy hypernetworks for controllable multi-objective re-ranking in recommender systems. By leveraging hypernetworks to generate parameters that balance multiple objectives, such as relevance and diversity, the system can dynamically adjust its recommendations based on user preferences and system goals. Similarly, HyperBandit  [12] maintains a neural network capable of generating the parameters for estimating time-varying rewards, taking into account the correlation between time features and user preferences. This approach enables dynamic adaptation to evolving user preferences over time. Mahabadi et al. [84] discuss parameter-efficient multi-task fine-tuning for transformers via shared hypernetworks, allowing for the efficient sharing of parameters across multiple tasks, reducing the overall computational load and improving the scalability of transformer models.
In conclusion, hypernetworks represent a powerful tool for parameter control in information retrieval and recommender systems. By enabling dynamic and task-specific parameter generation, hypernetworks can significantly improve the performance and adaptability of AI models, paving the way for more sophisticated and responsive systems.